OpenStudy (anonymous):

Rationalize the denominator of 3 over quantity of 6 minus square root of 5

OpenStudy (anonymous):

3/(6 - √5) = 3(6 + √5)/(6 - √5)(6 + √5) = (18 + 3√5)/(36 - 5) = (18 + 3√5)/31 the answer is d. (quantity of 18 plus 3 square root of 5 all over 31